OpenStudy (anonymous):

1+cos(2x) = cot(x)sin(2x)

hartnn (hartnn):

1+cos 2x = 2cos^2 x so u have to now prove cot x sin2x = 2cos^2 x write cot x as cos x/sin x and sin 2x as 2 sin x cos x got this? can u go further?

OpenStudy (anonymous):

Thank you for your response. This is as far as I was able to make it. I was able to simplify the LS to 2(1-sin^2x), although i'm not sure if this is correct. It's the RS that confuses me the most (below is the best i can do) \[\frac{ \cos(x) }{ \sin(x)} * 2\sin(x)\cos(x) \]Any further help would be appreciated (I'm helping Gohan with his math homework, wouldn't want to make him angry :P )

hartnn (hartnn):

your LS=2(1-sin^2 x)=2cos^2 x because sin^2x + cos^2 x = 1 in RS , u can cancel out sin x from numerator and denominator,which directly gives u 2 cos^2 x so LS=RS=2cos^2 x !! Hence,1+cos(2x) = cot(x)sin(2x)

OpenStudy (anonymous):

Wow, thank you! I did not know that you could cancel out the sin(x). Thank you for teaching me that, it will definitely help me in the future.

hartnn (hartnn):

welcome :)